### v1

@@ 4.2 Retargeting for RL Policy Training
RL Policy Training. For unstable systems like humanoid robots, the generated trajectory is not directly executable on the real robot due to the lack of feedback. One common solution is to train a RL policy to track the generated trajectory (He et al., 2024; Yang et al., 2025a; Liao et al., 2025). However, starting from infeasible human motion, general tracking policy training relies on heavy regularization (Ze et al., 2025; He et al., 2025) and complex curriculum design (He et al., 2024; Li et al., 2025c) to handle the noisy and infeasible reference motion. On the other hand, Liao et al. (2025); Yang et al. (2025a, b) shows with careful retargeting, the policy learning can be easier with less artifical designs. With trajectory retargeted by SPIDER, which already comes with a feedforward nominal control and a feasible trajectory, the RL only needs to learn a residual feedback term to correct the deviation from the nominal control:
@@
Training Results on Humanoid Robots. We evaluate the policy training with the adopted framework from Weng et al. (2025). We take standard practice of training PPO but removing the auxiliary contact reward for simplicity. Figure˜10 shows training progress on the OMOMO dataset along side with generated motion. The original human motion is offsetted from the object position, thus missing the contact. When training with the original human motion, the robot fails to grasp the object and only achieves body tracking goal. On the other hand, the policy trained with SPIDER is not only converged faster, but also achieves better object tracking performance. This again highlights the importance of physics grounding for reference motion.
@@ 5.1 Learning Manipulation from Human Data
Recent work explores learning robot skills from large-scale human videos by first detecting actions and transferring them to robots. One common strategy is kinematic retargeting, where human poses or keypoints are mapped to robot motions, as in DemoDiffusion (Park et al., 2025), OKAMI (Li et al., 2024), R+X (Papagiannis et al., 2025), and EgoZero (Liu et al., 2025). Another strategy is to train a human-centric policy and then adapt it to robots through fine-tuning with in-domain data, as in MimicPlay (Wang et al., 2023), Track2Act (Bharadhwaj et al., 2024), and VideoDex (Shaw et al., 2022). Our approach is complementary to these pipelines. Specifically, SPIDER can serve as a drop-in replacement for the human-to-robot action transfer components, providing more accurate and contact-aware mappings for dexterous and whole-body control.
@@ 5.2 Retargeting from Human Data
Kinematic Retargeting. These methods map human motion to robot configurations (Qin et al., 2022). While efficient and easy to compute, they often rely on specialized hardware (Xu et al., 2025), handcrafted motion primitives (Wu et al., 2024), and struggle with realism in contact-rich tasks (Qin et al., 2023). Being kinematic only, the generated motions are not compliant with physics constraints.
Learning-Based Retargeting Networks. Neural mapping approaches train networks to convert human motions into robot motions (Park et al., 2025; Yin et al., 2025b). Such models can outperform direct kinematic mappings and retain fast inference, but they require extensive pretraining and may fail when facing out-of-distribution motions or novel embodiments.
Optimization-Based Retargeting. Optimization-based approaches explicitly incorporate physics and contact constraints to ensure dynamical feasibility (Reda et al., 2023). They can generate high-quality, physically plausible motions, but often depend on detailed contact data (Lakshmipathy et al., 2025), specific data pipelines (Yang et al., 2025b), and strong priors (Nakaoka et al., 2005). Due to the non-convex natural of the problem, sampling-based approaches (Yang et al., 2025b) have emerged as a promising solution.
RL-Based Retargeting. RL has been used to retarget human demonstrations across embodiments (Lum et al., 2025; Li et al., 2025c). When combined with curriculum learning (Mandi et al., 2025; Liu et al., 2024), RL can produce dexterous, physically feasible robot motions. However, it typically requires training on each trajectory and significant computation, which limits scalability to internet-scale data and real-time deployment.
@@ 5.3 Sampling-based Optimization for Robot Control
Sampling-based optimization methods such as the cross-entropy method (De Boer et al., 2005), evolutionary algorithms (Salimans et al., 2017), and Bayesian optimization (Frazier, 2018) are powerful tools for solving non-convex and non-smooth problems. Due to their parallelizability and flexibility, these methods have been applied to navigation (Williams et al., 2016), legged locomotion (Xue et al., 2025), and dexterous manipulation (Howell et al., 2022; Li et al., 2025a). Despite their success in contact-rich control, they can suffer from instability and solution ambiguity in trajectory sampling (Kim et al., 2022). SPIDER addresses these challenges by guiding sampling with contact information, which helps preserve the intended contact sequence.

### v2

@@ 4.2 Retargeting for RL Policy Training
RL Policy Training. For unstable systems like humanoid robots, the generated trajectory is not directly executable on the real robot due to the lack of feedback. One common solution is to train a RL policy to track the generated trajectory (heOmniH2OUniversalDexterous2024; yang2025omniretarget; liao2025beyondmimic). However, starting from infeasible human motion, general tracking policy training relies on heavy regularization (zeTWISTTeleoperatedWholeBody2025; heASAPAligningSimulation2025) and complex curriculum design (heOmniH2OUniversalDexterous2024; liManipTransEfficientDexterous2025) to handle the noisy and infeasible reference motion. On the other hand, liao2025beyondmimic; yang2025omniretarget; yangPhysicsDrivenDataGeneration2025 shows with careful retargeting, the policy learning can be easier with less artifical designs. With trajectory retargeted by SPIDER, which already comes with a feedforward nominal control and a feasible trajectory, the RL policy only needs to learn a residual feedback term to correct the deviation from the nominal control:
@@
Training Results on Humanoid Robots. We evaluate the policy training with the adopted framework from weng2025hdmilearninginteractivehumanoid. We take standard practice of training PPO but removing the auxiliary contact reward for simplicity. Figure˜10 shows training progress on the OMOMO dataset along side with generated motion. The original human motion is offsetted from the object position, thus missing the contact. When training with the original human motion, the robot fails to grasp the object and only achieves body tracking goal. On the other hand, the policy trained with SPIDER is not only converged faster, but also achieves better object tracking performance. This again highlights the importance of physics grounding for reference motion.
@@ 5.1 Learning Manipulation from Human Data
Recent work explores learning robot skills from large-scale human videos by first detecting actions and transferring them to robots. One common strategy is kinematic retargeting, where human poses or keypoints are mapped to robot motions, as in DemoDiffusion (parkDemoDiffusionOneShotHuman2025), OKAMI (liOKAMITeachingHumanoid2024), R+X (papagiannisR+XRetrievalExecution2025), and EgoZero (liuEgoZeroRobotLearning2025). Another strategy is to train a human-centric policy and then adapt it to robots through fine-tuning with in-domain data, as in MimicPlay (wangMimicPlayLongHorizonImitation2023), Track2Act (bharadhwaj2024track2act), and VideoDex (shawVideoDexLearningDexterity2022). Our approach is complementary to these pipelines. Specifically, SPIDER can serve as a drop-in replacement for the human-to-robot action transfer components, providing more accurate and contact-aware mappings for dexterous and whole-body control.
@@ 5.2 Retargeting from Human Data
Kinematic Retargeting. These methods map human motion to robot configurations (qinOneHandMultiple2023). While efficient and easy to compute, they often rely on specialized hardware (xuDexUMIUsingHuman2025), handcrafted motion primitives (wuOneShotTransferLongHorizon2024), and struggle with realism in contact-rich tasks (qinAnyTeleopGeneralVisionBased2024). Being kinematic only, the generated motions are not compliant with physics constraints.
Learning-Based Retargeting Networks. Neural mapping approaches train networks to convert human motions into robot motions (parkDemoDiffusionOneShotHuman2025; yinDexterityGenFoundationController2025). Such models can outperform direct kinematic mappings and retain fast inference, but they require extensive pretraining and may fail when facing out-of-distribution motions or novel embodiments.
Optimization-Based Retargeting. Optimization-based approaches explicitly incorporate physics and contact constraints to ensure dynamical feasibility (redaPhysicsbasedMotionRetargeting2023). They can generate high-quality, physically plausible motions, but often depend on detailed contact data (lakshmipathyKinematicMotionRetargeting2024), specific data pipelines (yangPhysicsDrivenDataGeneration2025), and strong priors (nakaokaTaskModelLower2005). Due to the non-convex natural of the problem, sampling-based approaches (yangPhysicsDrivenDataGeneration2025; si2025exostartefficientlearningdexterous) have emerged as a promising solution.
RL-Based Retargeting. RL has been used to retarget human demonstrations across embodiments (lumCrossingHumanRobotEmbodiment2025; liManipTransEfficientDexterous2025). When combined with curriculum learning (mandiDexMachinaFunctionalRetargeting2025; liuQuasiSimParameterizedQuasiPhysical2024), RL can produce dexterous, physically feasible robot motions. However, it typically requires training on each trajectory and significant computation, which limits scalability to internet-scale data and real-time deployment.
@@ 5.3 Sampling-based Optimization for Robot Control
Sampling-based optimization methods such as the cross-entropy method (deboerTutorialCrossEntropyMethod2005), evolutionary algorithms (salimansEvolutionStrategiesScalable2017), and Bayesian optimization (frazierTutorialBayesianOptimization2018a) are powerful tools for solving non-convex and non-smooth problems. Due to their parallelizability and flexibility, these methods have been applied to navigation (williamsAggressiveDrivingModel2016), legged locomotion (xueFullOrderSamplingBasedMPC2024), and dexterous manipulation (howellPredictiveSamplingRealtime2022; liDROPDexterousReorientation2024). Despite their success in contact-rich control, they can suffer from instability and solution ambiguity in trajectory sampling (kimSmoothModelPredictive2022). SPIDER addresses these challenges by guiding sampling with contact information, which helps preserve the intended contact sequence.